OpenStudy (anonymous):

Simplify the complex fraction: @kx2bay

OpenStudy (anonymous):

\[-1\frac{ 36 }{ y^2 }\]\[\div 1-\frac{ 4 }{ y }-\frac{ 12 }{ y^2 }\]

OpenStudy (anonymous):

cookie u must hate fractions, dont u ;)

OpenStudy (anonymous):

I hate math period lol

OpenStudy (anonymous):

if I can interpret what you have above correctly, the answer will be  am a terrible drawer...